OpenStudy (anonymous):

Serena bought 3/4 pound of cheese. She used 2/3 of the cheese for sandwiches. How much cheese did Serena use for sandwiches? a- 1/4 b- 1/3 c- 1/2 d-5/7

OpenStudy (shamil98):

Find a common denominator first. In this case the common factor of 3 and 4 is 12. Multiply (3/4) by 3/3 and (-2/3) by 4/4 9/12 - 8/12 = 1/12

OpenStudy (anonymous):

2/3 of 3/4 ("of " means multiply) 2/3 * 3/4 = 6/12 = 1/2